# YOLO-based Bearing Fault Diagnosis With Continuous Wavelet Transform

### v2

@@ Abstract
This letter proposes a YOLO-based framework for spatial bearing fault diagnosis using time-frequency spectrograms derived from continuous wavelet transform (CWT). One-dimensional vibration signals are first transformed into time-frequency spectrograms using Morlet wavelets to capture transient fault signatures. These spectrograms are then processed by YOLOv9, v10, and v11 models to classify fault types. Evaluated on three benchmark datasets, including Case Western Reserve University (CWRU), Paderborn University (PU), and Intelligent Maintenance System (IMS), the proposed CWT-YOLO pipeline achieves significantly higher accuracy and generalizability than the baseline MCNN-LSTM model. Notably, YOLOv11 reaches mAP scores of 99.4% (CWRU), 97.8% (PU), and 99.5% (IMS). In addition, its region-aware detection mechanism enables direct visualization of fault locations in spectrograms, offering a practical solution for condition monitoring in rotating machinery.
@@ I Introduction
Rolling bearings are critical components in rotating machinery, and their operational reliability directly influences equipment lifespan, production efficiency, and safety [1]. Studies suggest that nearly 40% of rotating machinery failures originate from bearing faults [2]. These faults are often implicit in dynamic, non-stationary vibration signals, making their early-stage diagnosis particularly challenging. Vibration signals acquired from rotating machinery are typically recorded as one-dimensional time-domain sequences. These raw signals contain temporal patterns related to mechanical faults, such as impulsive features caused by inner race or ball defects [3]. However, these characteristics are often masked by noise and operational variability, especially during early-stage degradation [1]. Moreover, time-domain signals lack explicit frequency information, making it difficult to distinguish fault types that manifest at similar time intervals but differ in spectral behavior [2] (e.g., ball vs. outer race faults, inner race faults vs. imbalance, and looseness vs. misalignment, etc).
@@
To address these limitations, we propose a framework that combines CWT-based time-frequency spectrograms with You Only Look Once (YOLO)-based object detection models, including YOLOv9 [11], YOLOv10 [12], and YOLOv11 [13] to localize and classify fault patterns simultaneously. This approach leverages YOLO’s unified detection pipeline, which is optimized for both accuracy and speed, to enables fast and spatially explainable bearing fault diagnostics. To the best of our knowledge, this is the first study to explore YOLO-based spatial detection in time-frequency spectrograms for industrial bearing fault applications. The main contributions of this letter are as follows:
  • We propose a novel bearing fault diagnosis framework that combines CWT with YOLOv9, v10, and v11 to detect and classify fault patterns directly on time-frequency spectrograms.
  • The diagnostic task is reformulated as a spatial object detection problem, enabling region-aware visualization of fault patterns across time and frequency axes.
  • The proposed method is validated on three publicly available benchmark datasets, including Case Western Reserve University (CWRU) [14], Paderborn University (PU) [15], and the Intelligent Maintenance Systems (IMS) dataset [16], covering a wide range of operating conditions and fault types, and demonstrating strong generalizability.
  • This work is the first unified application of YOLOv9, v10, and v11 for spectrogram-based fault diagnosis, offering a transferable solution that balances high detection accuracy with model compactness and visual interpretability.
@@ II System Overview and Proposed CWT-YOLO
The overall architecture of the proposed bearing fault diagnosis system consists of three main stages: (1) time-frequency transformation using CWT, (2) spectrogram pre-processing and labeling, and (3) fault diagnosis using YOLO-based object detectors. Vibration signals acquired from rotating machinery are typically recorded as one-dimensional time-domain sequences. These raw signals contain temporal patterns related to mechanical faults, such as impulsive features caused by inner race or ball defects [3]. However, these characteristics are often masked by noise and operational variability, especially during early-stage degradation [1]. Moreover, time-domain signals lack explicit frequency information, making it difficult to distinguish fault types that manifest at similar time intervals but differ in spectral behavior [2], such as ball vs. outer race defects, early-stage inner race faults vs. imbalance, or looseness vs. misalignment, etc. Fig. 1 illustrates representative examples of such raw vibration signals. While some signal segments contain distinctive fluctuations, they are difficult to interpret visually or distinguish across fault types.
Figure 1: Sample vibration signals for four bearing conditions: (a) Normal, (b) Ball fault, (c) Inner race fault, (d) Outer race fault. Time (x-axis) and amplitude (y-axis) are shown for illustration.
@@ II-A Continuous Wavelet Transform (CWT)
In this work, we adopt the Morlet wavelet due to its favorable time-frequency localization and robustness to high-frequency noise. The Morlet wavelet combines a sinusoidal carrier with a Gaussian envelope $\psi(t)=e^{-\frac{t^{2}}{2}}\cos(5t)$, which offers balanced resolution across different frequency bands and has shown superior performance in early fault detection [7]. By applying CWT to raw vibration signals, we generate two-dimensional time-frequency spectrograms that expose transient features and scale-dependent energy distributions. These spectrograms serve as spatially structured inputs for subsequent object detection models.
@@ II-C2 Bounding Box Annotation for Object Detection
These annotations are saved in YOLO format, which specifies each object as a row of normalized values [class_id, x_center, y_center, width, height] with respect to the image dimensions. For consistency, each dataset is annotated separately to ensure accurate spatial alignment with the visual fault features.
@@ III-B Experimental Setup
All models are trained and evaluated on three public bearing fault datasets: CWRU [14], PU [15], and IMS [16]. The CWT spectrograms are generated from raw vibration signals with 2048-sample windows and $50\%$ overlap. Images are resized to $640\times 640$ pixels and annotated with bounding boxes using the LabelImg tool. The datasets are partitioned into 80% training, $10\%$ validation, and $10\%$ testing splits. All YOLO models are trained for $500$ epochs using stochastic gradient descent (SGD) (learning rate=0.01, batch size=8) on a workstation with Intel i7 CPU and RTX 3070Ti GPU.
@@ III-C Performance Comparison Across Datasets
Table I presents the mean average precision (mAP), precision (PRE), recall (REC), and F1 score (F1) for each model on the three datasets. All YOLO variants outperform the baseline MCNN-LSTM model [9] across all metrics and datasets. YOLOv11 consistently delivers the highest performance while also requiring the least inference time. The superior performance of YOLOv11 on the PU dataset is attributed to its spatial-attention mechanism (C2PSA), which effectively highlights localized patterns under varying load conditions. To evaluate the computational efficiency of each model, Table II demonstrates their theoretical FLOPs and parameter counts (Params). Values for the YOLO variants are computed using PyTorch on the same hardware setup described in Section III-B. For MCNN-LSTM, the values correspond to its original compact architecture designed for 1D time-series inputs without spectrogram transformation.
@@
Compared to YOLOv9, YOLOv11 reduces FLOPs by over 97% and parameters by nearly 95%, while achieving comparable or better accuracy across all datasets. Although MCNN-LSTM remains the most compact model in terms of complexity, its classification accuracy and generalizability are significantly lower. In contrast, YOLOv11 provides a balanced trade-off between efficiency and high detection accuracy, making it suitable for lightweight industrial fault monitoring. The proposed CWT-YOLO framework demonstrates strong generalization across diverse datasets. On the PU dataset, which includes multiple fault types under varying load and speed, YOLOv11 outperforms the baseline by over 16 percentage points in mAP. On the IMS dataset, all YOLO models achieve near-perfect classification, highlighting their ability to capture subtle degradation over long-term operation. Among the YOLO variants, YOLOv11 achieves the best overall performance, combining lower model complexity with faster inference than YOLOv9 and YOLOv10. These results indicate that the spatial detection paradigm, when applied to time-frequency spectrograms, supports accurate and interpretable fault localization.
@@ IV Conclusion
This letter presents a bearing fault diagnosis framework that combines CWT with YOLO-based object detection to leverage time-frequency spectrograms for localized classification. By converting one-dimensional vibration signals into spatially structured spectrograms, the system enables visual identification of fault patterns using bounding box detection. Evaluations on the CWRU, PU, and IMS datasets confirm that YOLO-based models significantly outperform the MCNN-LSTM baseline in classification accuracy, with YOLOv11 achieving the best balance between detection performance and computational complexity. These results highlight the potential of spectrogram-level object detection for building practical, interpretable, and deployment-ready condition systems.

### v3

@@ Abstract
This letter presents a locality-aware bearing fault diagnosis framework that operates on time-frequency representations and enables spatially interpretable decision-making. One-dimensional vibration signals are first mapped to two-dimensional time-frequency spectrograms using the continuous wavelet transform (CWT) with Morlet wavelets to enhance transient fault signatures. The diagnosis task is then formulated as object detection on the time-frequency plane, where YOLOv9, YOLOv10, and YOLOv11 are employed to localize fault-relevant regions and classify fault types simultaneously. Experiments on three public benchmarks, including Case Western Reserve University (CWRU), Paderborn University (PU), and Intelligent Maintenance System (IMS), demonstrate strong cross-dataset generalization compared with a representative MCNN-LSTM baseline. In particular, YOLOv11 achieves mAP@0.5 of 99.0% (CWRU), 97.8% (PU), and 99.5% (IMS), while providing region-aware visualization of fault patterns in the time-frequency domain. These results suggest that detection-based inference on CWT spectrograms provides an effective and interpretable complementary approach to conventional global classification for rotating machinery condition monitoring.
@@ I Introduction
Rolling bearings are critical components in rotating machinery, and their operational reliability directly influences equipment lifespan, production efficiency, and safety [1]. Prior studies report that nearly 40% of rotating machinery failures originate from bearing faults [2]. These faults are often embedded in dynamic, non-stationary vibration signals, making their early-stage diagnosis particularly challenging. Vibration signals acquired from rotating machinery are typically recorded as one-dimensional time-domain sequences. These raw signals contain temporal patterns related to mechanical faults, such as impulsive features caused by inner-race or ball defects [3]. However, these characteristics are often masked by noise and operational variability, especially during early-stage degradation [1]. Moreover, time-domain signals lack explicit frequency information, making it difficult to distinguish fault types that manifest at similar time intervals but differ in spectral behavior [2] (e.g., ball vs. outer-race faults, inner-race faults vs. imbalance, and looseness vs. misalignment).
@@
To address these limitations, we propose a detection-based diagnosis framework that explicitly exploits locality in the time-frequency plane. Specifically, we first apply Morlet-based CWT to convert vibration segments into time-frequency spectrograms, where fault-related transients manifest as spatially localized energy patterns. We then cast bearing diagnosis as an object detection problem on these spectrograms and adopt YOLO-based detectors, including YOLOv9 [11], YOLOv10 [12], and YOLOv11 [13], to jointly localize fault-relevant regions and classify fault types. Compared with conventional global classification, the proposed formulation enables region-aware interpretability and can be more robust to operating variations because decisions are made based on localized time-frequency evidence. To the best of our knowledge, this is the first study to systematically investigate modern YOLO variants (v9–v11) for spectrogram-level fault localization and diagnosis across multiple public bearing benchmarks.
The main contributions of this letter are as follows:
  • We introduce a locality-aware diagnosis pipeline that combines Morlet-based CWT with detection-based inference on time-frequency representations for bearing fault analysis.
  • We reformulate bearing fault diagnosis from global segment classification to object detection on the time-frequency plane, enabling spatially interpretable localization of fault-relevant regions.
  • We provide an extensive evaluation on three public benchmarks, including CWRU [14], PU [15], and IMS [16], demonstrating strong generalization under diverse operating conditions and fault types.
  • We benchmark three modern YOLO variants (YOLOv9–YOLOv11) under a unified spectrogram-based detection setting, highlighting the accuracy and efficiency trade-off for deployment-oriented condition monitoring.
@@ II System Overview and Proposed CWT-YOLO
The proposed system consists of three stages: (1) CWT-based time-frequency transformation, (2) spectrogram pre-processing and bounding-box annotation, and (3) YOLO-based detection for joint fault localization and classification. Fig. 1 shows representative vibration segments under different bearing conditions, where discriminative transients are often difficult to identify directly in the time domain.
Figure 1: Sample vibration signals for four bearing conditions: (a) Normal, (b) Ball fault, (c) Inner race fault, (d) Outer race fault. Time (x-axis) and amplitude (y-axis) are shown for illustration, motivating a time-frequency transformation to expose discriminative fault signatures.
@@ II-A Continuous Wavelet Transform (CWT)
In this work, we adopt the Morlet wavelet due to its favorable time-frequency localization and robustness to high-frequency noise. The Morlet wavelet combines a sinusoidal carrier with a Gaussian envelope, providing favorable time-frequency localization and robustness to high-frequency noise. In our implementation, we use a real-valued Morlet form $\psi(t)=e^{-\frac{t^{2}}{2}}\cos(5t)$, which offers balanced resolution across different frequency bands and has shown strong performance for early fault detection [7]. By applying CWT to raw vibration signals, we generate two-dimensional time-frequency spectrograms that expose transient features and scale-dependent energy distributions. These spectrograms serve as spatially structured inputs for subsequent object detection models.
@@ II-C2 Bounding Box Annotation for Object Detection
These annotations are saved in YOLO format, which specifies each object as a row of normalized values [class_id, x_center, y_center, width, height] with respect to the image dimensions. For consistency, each dataset is annotated separately to ensure accurate spatial alignment with the visual fault features. To improve reproducibility, we adopt a consistent annotation criterion across datasets. Bounding boxes are drawn to cover the most prominent time-frequency energy concentrations associated with each fault class, while excluding background regions with diffuse responses. When multiple transient blobs appear in a spectrogram, the box encloses the dominant cluster that is visually persistent across samples of the same class, which reduces subjective variation in ROI selection.
@@ III-B Experimental Setup
All models are trained and evaluated on three public bearing fault datasets: CWRU [14], PU [15], and IMS [16]. The CWT spectrograms are generated from raw vibration signals with 2048-sample windows and $50\%$ overlap. Images are resized to $640\times 640$ pixels and annotated with bounding boxes using the LabelImg tool. The datasets are partitioned into 80% training, $10\%$ validation, and $10\%$ testing splits. All YOLO models are trained for $500$ epochs using stochastic gradient descent (SGD) (learning rate=0.01, batch size=8) on a workstation with an Intel Core i7-class CPU and an NVIDIA RTX 3070 Ti GPU. Unless otherwise stated, the same training protocol and data split are used for all YOLO variants to ensure a fair comparison.
@@ III-C Performance Comparison Across Datasets
Table I reports the mean average precision (mAP@0.5), precision (PRE), recall (REC), and F1 score (F1) for each model on the three datasets. Overall, YOLO-based detectors consistently outperform the MCNN-LSTM baseline [9] across datasets, indicating that detection-based inference on time-frequency representations can better capture localized fault signatures. Notably, YOLOv11 achieves the strongest results on PU and competitive performance on CWRU and IMS, suggesting a favorable accuracy–efficiency trade-off among the evaluated variants. The strong PU performance of YOLOv11 can be partly explained by its attention mechanism (C2PSA), which helps emphasize localized patterns under varying operating conditions.
To compare computational complexity, Table II lists the estimated FLOPs and parameter counts (Params) under the adopted implementations and the same input resolution. Values for YOLO variants are computed using PyTorch on the hardware described in Section III-B. For MCNN-LSTM, the values correspond to its original compact 1D time-series architecture.
@@
Compared with YOLOv9, YOLOv10 and YOLOv11 substantially reduce computational complexity, while maintaining competitive detection performance across datasets. Although YOLOv9 achieves slightly higher mAP on CWRU, YOLOv11 provides a favorable balance between accuracy and efficiency, and it performs particularly well on PU and IMS. These results indicate that detection-based inference on CWT spectrograms can generalize across datasets with diverse operating conditions, while enabling interpretable fault localization through region-aware predictions.
It is also worth noting that the complexity values in Table II are reported for reference under the adopted implementations and input resolution, and the complexity between 2D spectrogram-based detectors and 1D time-series classifiers is not directly comparable. Nevertheless, the proposed framework suggests a practical accuracy and interpretability gain when leveraging localized time-frequency evidence for bearing fault diagnosis.
@@ IV Conclusion
This letter presents a locality-aware bearing fault diagnosis framework that combines Morlet-based CWT with YOLO-based object detection on time-frequency spectrograms. By casting diagnosis as detection on the time-frequency plane, the proposed approach enables localization-driven inference and provides region-level interpretability by highlighting fault-relevant patterns. Experiments on the CWRU, PU, and IMS datasets demonstrate that YOLO-based detectors substantially improve generalization compared with the MCNN-LSTM baseline, while YOLOv11 offers a favorable balance between detection performance and computational complexity. These findings suggest that spectrogram-level detection is a practical and interpretable complementary approach to conventional global classification for rotating machinery condition monitoring.